# Preference-Conditioned Gradient Variations for Multi-Objective Quality-Diversity

### v1

@@ V-C Metrics
  1. 1.
    The moqd-score [21, 25] is the sum of the hypervolumes of the Pareto fronts stored in the archive $\mathcal{A}$:
@@
    This metric aims to assess if an algorithm can find high-performing Pareto fronts, for a range of features.
  2. 2.
    We introduce the moqd-sparsity-score, which we define as the average sparsity of each Pareto front $\Xi(\mathcal{P}_{i})$ of the archive:
@@
  3. 3.
    The global-hypervolume is the hypervolume of the Pareto front formed over all of the solutions in the archive (which we term the global Pareto front). The metric assesses the elitist performance of an algorithm. That is, the performance of solutions on the objective functions that are possible when disregarding the solution’s feature.
  4. 4.
    By the same reasoning as the moqd-sparsity-score, we also introduce the global-sparsity, which is the sparsity of the Pareto front formed over all of the solutions in the archive.
  5. 5.
@@
  6. 6.
    The coverage is the proportion of cells in of an archive that are occupied. It reflects how many different features the algorithm is able to uncover (regardless of the performance of the solutions). Since all of the algorithms achieved a similar performance on this metric, we report the results in the supplementary materials.
@@
V-D Hyperparameters
Figure 4: moqd-score, global-hypervolume and maximum sum of scores (Section V-C) for mome-p2c compared to all baselines across all tasks. Each experiment is replicated 20 times with random seeds. The solid line is the median and the shaded area represents the first and third quartiles.
@@ V-D Hyperparameters
All experiments were run for the same total budget of $4000$ iterations with a batch size of 256 evaluations per generation, corresponding to a total of $1,024,000$ evaluations. We used CVT tessellation [31] to create archives with $128$ cells, each with a maximum of Pareto Front length of $50$. For all experiments, we use a Iso+LineDD operator [51] as the ga variation operator, with $\sigma_{1}=0.005$ and $\sigma_{2}=0.05$. The reference points for each environment and the actor-critic training parameters were kept the same across all algorithms and are provided in the supplementary material.
VI Results
In this section, we present the results for all baselines. Each experiment is replicated 20 times with random seeds. We report $p$-values based on the Wilcoxon–Mann–Whitney $U$ test with Holm-Bonferroni correction to ensure statistical validation of the results [52, 53].
Figure 5: Boxplots to display sparsity metrics calculted on the final archive of mome-p2c and mome-pgx over 20 replications. The labels A2, A3, HC2, H2, H3 and W2 correspond to the Ant-2, Ant-3, HalfCheetah-2, Hopper-2, Hopper-3 and Walker-2 environments respectively.
@@ VI Results
Figure 6: moqd-score (Section V-C) for mome-p2c compared to all ablations across all tasks. Each experiment is replicated 20 times with random seeds. The solid line is the median and the shaded area represents the first and third quartiles.
VI-A Main Results
@@
The experimental results presented in Figure 4 demonstrate that mome-p2c outperforms or matches all baselines on all tasks and all metrics. mome-p2c achieves a significantly higher moqd-score than all baselines on Ant-2, Ant-3, HalfCheetah-2 and Hopper-3 ($p<0.01$). mome-p2c matches the moqd-score of mome-pgx, the previous state-of-the-art, on the remaining environments Hopper-2 and Walker-2, but at a lower storage and computational cost. Crucially, in scenarios where mome-p2c does not markedly outperform mome-pgx, it still attains lower sparsity scores (Figure 5), indicating that it achieves smoother array of trade-offs.
@@
mome-p2c also outperforms nsga-ii and spea2 on the global-hypervolume metric, algorithms that specifically aim to maximise this metric. Furthermore, mome-p2c achieves a better maximum sum of scores than pga-me across all tasks ($p<10^{-5}$) except Hopper-2 and Walker-2 where it still shows improved but not statistically significant performance. Importantly, we see that mome-p2c achieves better performance on tri-dimensional tasks, affirming not only its computational efficiency but also its scalability in handling more complex tasks with an increased number of objectives.
VI-B Ablations
VI-B1 Ablation studies
@@
In our ablation studies, we evaluate mome-p2c against mome-pgx together with four distinct modifications to understand the contribution of each component in mome-p2c. These ablations include:
@@
VI-B2 Results
@@
The results from our ablation studies provide a deeper understanding of the individual components contributing to mome-p2c’s effectiveness. Notably, the no-actor ablation, which removes the actor injection mechanism from mome-p2c, shows an interesting pattern. In the Ant-2 and HalfCheetah-2 ($p<10^{-4}$), mome-p2c significantly outperforms the no-actor ablation, suggesting that some tasks which require a more nuanced exploration strategy may benefit greatly from the the actor injection mechanism. Moreover, in all of the other tasks except for Hopper-2, no-actor either matches or falls behind the full mome-p2c model.
Furthermore, the no-crowding ablation, where mome-p2c operates without its crowding mechanisms, significantly underperforms compared to the standard mome-p2c across all tasks ($p<10^{-4}$). This uniform selection and replacement strategy evidently lack the refined search capabilities provided by crowding, underscoring the importance of these mechanisms in guiding the algorithm towards more diverse and high-quality solutions.
Finally, the modifications in preference sampling strategies, as explored in the keep-pref and one-hot ablations, markedly influence mome-p2c ’s performance. The keep-pref ablation, which retains the parent’s preference in the policy-gradient variation operator shows no improvement over mome-p2c across all tasks. The one-hot ablation, employing one-hot preferences in equal batch sizes achieves equal or worse moqd-score compared to mome-p2c across all tasks except Hopper-2. These results highlight the critical role of preference management within mome-p2c and open up a promising avenue for future research. Specifically, developing strategies to predict preferences that might lead to significant hypervolume improvements, could further enhance mome-p2c’s performance, especially in complex multi-objective optimization tasks [38].
VII Conclusion
@@
In this paper, we have introduced a novel algorithm, mome-p2c, which represents a significant advancement in the field of Multi-Objective Quality-Diversity (MOQD) optimization. Our experiments and ablation studies have demonstrated mome-p2c’s ability to balance multiple objectives effectively, outperforming existing state-of-the-art methods in challenging continuous control environments.
@@
One of the key strengths of mome-p2c is its use of preference-conditioned policy gradient mutations, which not only enhance the exploration process but also ensures an even distribution of solutions across the Pareto front. This approach addresses the limitations of mome-pgx that struggled with scalability and efficiency. Furthermore, mome-p2c’s ability to perform well in tri-dimensional tasks highlights its scalability and adaptability to more complex and realistic scenarios. Our ablation studies highlight that the strategy for sampling preferences can have a large impact on the performance of mome-p2c. The exploration of using models to predict which preference will lead to the largest hypervolume gain [38] presents an exciting direction for further research.
@@ Appendix A Coverage Results
Figure 7: Median coverage performance of 20 seeds, the shaded regions show the inter-quartile range.
Figure 7 presents the coverage results for all of the baseline algorithms. As expected, all Quality-Diversity algorithms achieve a higher coverage score than the MOEA baselines, as they explicitly seek diverse solutions.
@@ Appendix B Hypervolume Reference Points
Table II presents the reference points used to calculate the hypervolume metrics in each of the tasks. The same reference points were used for all of the experiments.
TABLE II: Reference points
@@ Appendix C Policy Gradient Hyperparameters
Table III presents all of the policy gradient hyperparameters that are used for our algorithms. All hyperparameters were kept the same for each task and for all algorithms which used PG variations.
TABLE III: Policy Gradient Network Hyperparameters

### v2

@@ 5.3. Metrics
  1. (1)
    The moqd-score (Pierrot et al., 2022b; Janmohamed et al., 2023) is the sum of the hypervolumes of the non-dominated fronts stored in the archive $\mathcal{A}$:
@@
    This metric aims to assess if an algorithm can find high-performing fronts, for a range of features.
  2. (2)
    We introduce the moqd-sparsity-score, which we define as the average sparsity of each non-dominated front $\Xi(\mathcal{P}_{i})$ of the archive:
@@
  3. (3)
    The global-hypervolume is the hypervolume of the non-dominated front formed over all of the solutions in the archive (which we term the global non-dominated front). The metric assesses the elitist performance of an algorithm. That is, the performance of solutions on the objective functions that are possible when disregarding the solution’s feature.
  4. (4)
    By the same reasoning as the moqd-sparsity-score, we also introduce the global-sparsity, which is the sparsity of the non-dominated front formed over all of the solutions in the archive.
  5. (5)
@@
  6. (6)
    The coverage is the proportion of cells in of an archive that are occupied. It reflects how many different features the algorithm is able to uncover (regardless of the performance of the solutions). Since all of the algorithms achieved a similar performance on this metric, we report the results for baselines and ablations in Appendix A .
@@
5.4. Hyperparameters
@@
All experiments were run for the same total budget of of $1,024,000$ evaluations. In mome-p2c, mome-pgx, mome, nsga-ii and spea2 this corresponded to $4000$ iterations with a batch size of 256 evaluations per generation. In nsga-ii-small and spea2-small, we used a batch size of $1600$ which was the limit of our GPU and ran the algorithm for $640$ iterations. We used CVT tessellation (Vassiliades et al., 2018) to create archives with $128$ cells, each with a maximum of non-dominated Front length of $50$. For all experiments, we use a Iso+LineDD operator (Vassiliades and Mouret, 2018) as the ga variation operator, with $\sigma_{1}=0.005$ and $\sigma_{2}=0.05$. This operator mutates a solution by adding isotropic Gaussian noise scaled by $\sigma_{1}$ and a directional perturbation along the vector between two individuals, scaled by $\sigma_{2}$, promoting both exploration and guided variation. The reference points for each environment and the actor-critic training parameters were kept the same across all algorithms and are provided Appendix B and Appendix C.
@@
Figure 4. moqd-score, global-hypervolume and maximum sum of scores (Section 5.3) for mome-p2c compared to all baselines across all tasks. Each experiment is replicated 20 times with random seeds. The solid line is the median and the shaded area represents the first and third quartiles.
6. Results
In this section, we present the results for all baselines. Each experiment is replicated 20 times with random seeds. We report $p$-values based on the Wilcoxon–Mann–Whitney $U$ test with Holm-Bonferroni correction to ensure statistical validation of the results (Wilcoxon, 1992; Holm, 1979).
./Refer to caption
Figure 5. Boxplots to display sparsity metrics calculated on the final archive of mome-p2c and mome-pgx over 20 replications. The labels A2, A3, HC2, H2, H3 and W2 correspond to the ant-2, ant-3, halfcheetah-2, hopper-2, hopper-3 and walker-2 environments respectively.
Figure 6. moqd-score (Section 5.3) for mome-p2c compared to all ablations across all tasks. Each experiment is replicated 20 times with random seeds. The solid line is the median and the shaded area represents the first and third quartiles.
6.1. Main Results
@@
The experimental results presented in Figure 4 demonstrate that mome-p2c outperforms or matches all baselines on all tasks and all metrics. mome-p2c achieves a significantly higher moqd-score than all baselines on ant-2, ant-3, halfcheetah-2 and hopper-3 ($p<0.02$). mome-p2c matches the moqd-score of mome-pgx, the previous state-of-the-art, on the remaining environments hopper-2 and walker-2. We include archive visualisations for environments with two-dimensional features in Appendix E to further visualise these results. Crucially, in the tasks where mome-p2c does not markedly outperform mome-pgx on moqd-score, it still attains lower moqd-sparsity-score (Figure 5), indicating that it achieves smoother array of trade-offs for each feature.
mome-p2c also outperforms or matches all baselines on the global-hypervolume metric. We include visualisations of the global non-dominated Fronts obtained by representative runs of each of the algorithms in Appendix F. mome-p2c outperforms nsga-ii, nsga-ii-small, spea2 and spea2-small on the global-hypervolume in all environments, algorithms that specifically aim to maximise this metric. We hypothesise that this result is due to the policy-gradient variations of mome-p2c, which improve the search capabilities of these algorithms in high-dimensional search spaces with neural network controllers. However, mome-p2c still outperforms or matches the global-hypervolume of the other baselines which use policy-gradient variations (pga-me and mome-pgx), demonstrating the benefits of using preference-conditioned updates. Importantly, in the environments where mome-p2c matches mome-pgx on global-hypervolume (hopper-2, hopper-3, walker-2), it still achieves a lower global-sparsity as illustrated in Figure 5.
@@
Furthermore, mome-p2c achieves a better maximum sum of scores than pga-me across all tasks ($p<10^{-5}$) except hopper-2 and walker-2 where it still shows improved but not statistically significant performance. This highlights that even if a specific reward preference is available, leveraging multiple objectives may still result in better exploration and higher returns compared to single-objective baselines. Furthermore, we see that mome-p2c achieves better performance on tri-dimensional tasks, affirming not only its computational efficiency but also its scalability in handling more complex tasks with an increased number of objectives.
6.2. Ablations
6.2.1. Ablation studies
@@
In our ablation studies, we evaluate mome-p2c against mome-pgx together with five distinct modifications to understand the contribution of each component in mome-p2c. These ablations include:
@@
  • no-ga: mome-p2c without any genetic mutations. At each generation, there are 128 policy-gradient offspring and 128 actor-injection offspring.
6.2.2. Results
The results from our ablation studies provide a deeper understanding of the individual components contributing to mome-p2c’s effectiveness, we include further archive visualisations in Appendix E. Firstly, we examine the modifications in preference sampling strategies, as explored in the keep-pref and one-hot ablations. mome-p2c matches or outperforms the keep-pref ablation, which retains the parent’s preference in the policy-gradient variation operator, across all tasks except walker-2. Similarly, the one-hot ablation, employing one-hot preferences in equal batch sizes has no statistical difference or is worse compared to mome-p2c across all tasks in all metrics except walker-2. These results highlight that mome-p2c is robust to the preference-sampling strategy in most environments. However, the results on walker-2 suggests that, in some environments, alternative sampling strategies may be optimal. We leave this as a promising avenue for future research. Additionally, we note that one-hot outperforms or matches mome-pgx across all metrics, in all environments, highlighting that simply using the preference-conditioned critic leads to improved performance.
The no-crowding ablation, where mome-p2c operates without its crowding mechanisms, significantly underperforms compared to mome-p2c across all tasks ($p<10^{-4}$). This highlights that the crowding-based selection and addition mechanism is crucial for guiding the algorithm towards more diverse and high-quality solutions. Likewise, mome-p2c also outperforms the no-ga ablation in every environment except for ant-2, where the no-ga ablation environment achieves a higher global-hypervolume and maximum sum of scores. We hypothesise that since the Ant robot has higher dimensionality than other morphologies, in this environment the high-dimensionality of the search space requires more gradient-based search power. However, the no-ga has a worse performance across all other metrics in all of the other environments, particularly hopper-2, hopper-3 and walker-2, highlighting that in general the genetic variations are an essential component of the mome-p2c algorithm.
@@ 6.2.2. Results
Finally, we consider the no-actor ablation, which removes the actor injection mechanism from mome-p2c. In the ant-2 and halfcheetah-2 ($p<10^{-4}$), mome-p2c markedly outperforms the no-actor ablation, suggesting that the actor injection mechanism is essential in these tasks. Moreover, no-actor either matches or falls behind the full mome-p2c model in all other tasks and all other metrics, except for the moqd-score in ant-3.
7. Conclusion
@@
In this paper, we have introduced a novel algorithm, mome-p2c, which represents a significant advancement in the field of Multi-Objective Quality-Diversity (MOQD) optimisation. Our experiments and ablation studies have demonstrated mome-p2c’s ability to balance multiple objectives effectively, outperforming existing state-of-the-art methods in challenging continuous control environments.
@@
One of the key strengths of mome-p2c is its use of preference-conditioned policy gradient mutations, which not only enhance the exploration process but also ensures an even distribution of solutions across the non-dominated front. This approach addresses the limitations of mome-pgx that struggled with scalability and efficiency. Furthermore, mome-p2c’s ability to perform well in tri-dimensional tasks highlights its scalability and adaptability to more complex and realistic scenarios.
Beyond these results, a limitation of our study is that experiments were restricted to standard Brax locomotion benchmarks with up to three objectives; extending evaluation to higher-dimensional and non-robotics tasks remains an important avenue for future work. Additionally, a key limitation of mome-p2c lies in how preferences are currently sampled. While the algorithm is generally robust across environments, certain tasks, such as walker-2, demonstrate that naive or uniform preference sampling may be suboptimal. This suggests that dynamic or adaptive sampling strategies could further improve performance, especially in environments with complex or shifting trade-offs. The exploration of using models to predict which preference will lead to the largest hypervolume gain (Xu et al., 2020) presents an exciting direction for further research.
Finally, while our parameter comparisons in Appendix C demonstrate that mome-p2c is more memory-efficient than mome-pgx, scaling to higher-dimensional objective spaces introduces additional challenges. In particular, we expect random preference sampling to become increasingly ineffective as dimensionality grows and, as noted in Section 4.1, crowding distance is known to lose effectiveness in many-objective optimisation (Zheng et al., 2024; Köppen and Yoshida, 2007). Addressing these issues through adaptive preference sampling, scalable diversity-preservation mechanisms, and architectural innovations represents a key line of future work in extending the generality of mome-p2c.
Acknowledgements
This work was supported by PhD scholarship funding for Hannah from InstaDeep.
@@ Appendix A Coverage Results
Figure 7 and Figure 8 present the coverage results for all of the baseline and ablations algorithms respectively. As expected, all Quality-Diversity algorithms achieve a higher coverage score than the MOEA baselines, as they explicitly seek diverse solutions.
Figure 7. Median coverage performance of 20 seeds, the shaded regions show the inter-quartile range.
Figure 8. Median coverage performance of 20 seeds, the shaded regions show the inter-quartile range.
@@ Appendix B Hypervolume Reference Points
Table 2 presents the reference points used to calculate the hypervolume metrics in each of the tasks. The same reference points were used for all of the experiments.
Table 2. Reference points
@@ Appendix C Policy Gradient Hyperparameters
Table 3 presents all of the policy gradient hyperparameters that are used for our algorithms. All hyperparameters were kept the same for each task and for all algorithms which used PG variations.
Table 3. Policy Gradient Network Hyperparameters
@@
In addition to hyperparameters, we report the number of parameters required for actor and critic networks in mome-p2c compared to mome-pgx across all environments. These values directly support the argument in the main text regarding the improved memory efficiency of mome-p2c.
Table 4. Number of parameters in total for Critic and Actor Networks for mome-p2c and mome-pgx in each environment.
mome-p2c mome-pgx
ant-2 182,788 362,500
ant-3 183,814 543,750
halfcheetah-2 146,436 289,796
hopper-2 141,316 279,556
hopper-3 142,342 419,334
walker-2 145,924 288,772
(a) Critic Networks
mome-p2c mome-pgx
ant-2 90,376 180,752
ant-3 90,376 271,128
halfcheetah-2 72,198 144,396
hopper-2 69,635 139,270
hopper-3 69,635 208,905
walker-2 71,942 143,884
(b) Actor Networks
As shown, mome-p2c consistently requires roughly half the number of parameters compared to mome-pgx across all environments, confirming its memory efficiency advantage. This reduction holds for both actor and critic networks, and becomes more pronounced as the number of objectives increases (e.g., ant-3 and hopper-3). The difference in parameter counts between ant-2 and ant-3, and similarly between hopper-2 and hopper-3, illustrates how scaling to additional objectives affects the two approaches. For mome-pgx, the number of parameters grows roughly linearly with the number of objectives, since a separate actor–critic pair is maintained for each objective. In contrast, mome-p2c shows only a marginal increase in parameters between the two- and three-objective tasks, as the same networks are reused and only the input dimensionality grows. This highlights that while both methods scale with the number of objectives, mome-p2c ’s architecture is more memory-efficient and thus better suited to higher-dimensional problems.
Figure 9. Wall-clock training time (in minutes) for policy-gradient mutation steps and actor–critic training in each environment. mome-p2c consistently requires less time than mome-pgx across all tasks. Note: timings were collected from un-jitted runs for accurate measurement; actual JIT-compiled training is substantially faster.
To further support our claims of computational efficiency, Figure 9 reports the wall-clock time of policy-gradient mutation steps and actor–critic training for mome-p2c and mome-pgx across all environments. We find that mome-p2c consistently requires less training time than mome-pgx in every environment, demonstrating its improved computational efficiency. To obtain accurate timing measurements, we disabled JIT compilation in the QDax framework, as JIT fusion obscures intermediate operation timings. As a result, the reported values are significantly longer than those observed in practice, but they provide a fair relative comparison across methods. In typical JIT-compiled runs, each environment for mome-p2c completes in under one hour depending on the task. Because un-jitted runs are substantially slower, we report results for five random seeds per environment. Overall, the lower wall-clock times and reduced parameter counts together confirm that mome-p2c achieves both computational and memory efficiency over mome-pgx.
Appendix D Ablation – Effect of Archive Size for PGA
In the main text, we compared mome-p2c to pga-me using an archive size of $k\times|\mathcal{P}|$ cells, to to ensure a fair comparison in terms of archive capacity. However, having a large number of cells may decrease the selection pressure of the algorithm and therefore negatively impact the performance of pga-me. We therefore reran pga-me with the same number of grid cells $k$ as the MOQD methods. The results are reported in Figure 10.
Figure 10. Comparison between mome-p2c and pga-me when both use the same number of grid cells $k$. Each curve shows the mean and interquartile range over 20 runs. Results demonstrate that mome-p2c consistently outperforms or matches pga-me across environments, even when controlling for archive size.
We observe that mome-p2c continues to achieve a higher moqd-score across all tasks ($p<0.001$). mome-p2c also significantly outperforms pga-me on the maximum sum of scores metrics in most environments, except walker-2 and hopper-2 where results are not statistically significant. The same trend is observed for the global-hypervolume, where mome-p2c significantly outperforms pga-me on ant-2, ant-3, and halfcheetah-2 ($p<10^{-4}$), as well as hopper-3 ($p<0.01$). Taken together, these results confirm that the observed improvements are not merely an artefact of archive size: even when controlling for the number of cells, mome-p2c consistently outperforms or matches pga-me across environments.
Appendix E Repertoire Visualisations
To provide additional insight into the behaviour of the algorithms, we visualise the final repertoires in the feature space for environments with two-dimensional feature descriptors. Figures 11 and 12 show representative seeds for each baseline and ablation, respectively.
Figure 11. Representative repertoires for the main baselines in experiments with two-dimensional features. Colours indicate the hypervolume of the non-dominated set in each cell. mome-p2c achieves broader coverage and higher-quality solutions compared to other baselines.
In Figure 11, we observe that mome-p2c produces larger and denser archives with non-dominated fronts with higher hypervolume than the other baselines. Compared to population-based MOEAs such as nsga-ii and spea2, which often leave large regions of the feature space unexplored, mome-p2c and other moqd algorithms cover the space more uniformly. We also note that that the other baselines achieve good coverage, but fail to discover as high-performing fronts as mome-p2c.
Figure 12. Representative repertoires for the ablation studies in experiments with two-dimensional features. Each ablation shows a noticeable drop in coverage and/or quality compared to mome-p2c, consistent with the quantitative results in the main paper.
Figure 12 aligns with the quantitative results in Section 6.1, where we observe that in the halfcheetah-2 tasks mome-p2c performs much better than the ablations, but the ablations are competitive in walker-2 environment.
Appendix F Global Non-Dominated Fronts
Figure 13 and Figure 14 presents the visualisations of the global non-dominated fronts obtained by each of the baseline and ablation algorithms.
Figure 13. Visualisations of global non-dominated Fronts obtained by representative runs of each of the baseline algorithms.
Figure 14. Visualisations of global non-dominated Fronts obtained by representative runs of each of the ablation algorithms.